# Structure of activity in multiregion recurrent neural networks

### v1

@@ V.4 Dimension and topology of the attractor manifold
The solution set of this linear program, called the feasible region, is a convex polytope. The dimension $d$ of the feasible region is, barring fine tuning, the number of variables minus the number of constraints,
$d=n-M=\frac{M(M-3)}{2}.$ (34)
This is also the dimension of the current-space manifold. Thus, for $M\geq 4$, the manifold is continuous with dimension $d\geq 2$. For $M=3$, the manifold reduces to a zero-dimensional point set. For $M=2$, or for a sufficiently nonuniform constraint vector $\bm{A}$, the linear program is infeasible (i.e., has no solutions). In this case, the assumptions of the linear-program formulation are violated and we must return to the fixed-point equations. We revisit this momentarily.
To characterize the topology of the current-space manifold, we first observe that, for a given point $\bm{x}$ on the feasible region, there are $2^{n}$ corresponding points in current space. This multiplicity arises from the different ways one can choose the signs of the currents. The connectedness, or lack thereof, of current space hinges on whether $x_{k}=0$ for each $k\in\{1,\ldots,n\}$ is included in the feasible region. If included, positive and negative current-space branches connect; otherwise, a binary fracture of the manifold is induced. We refer to variables $x_{k}$ that never take on zero as fracture variables and denote their number by $f$. This is visualized in Fig. 6(a). Each fracture variable contributes one binary split to the manifold, resulting in $2^{f}$ connected components. Zeros in components of $\bm{x}$ occur only at vertices of the feasible region, so to identify all fracture variables, it suffices to enumerate all vertices. This can be done using the double-description method of Motzkin [51, 52]. Alternatively, to determine if $x_{k}$ is a fracture variable, one could set $x_{k}=0$ and check if the resulting linear program in $n-1$ variables is feasible using the efficient simplex algorithm.
To generate many realizations of this linear program, we repeatedly do the following. We first pick a vector $\bm{b}$ by sampling its components $b^{\mu}$ uniformly over $1\leq b^{\mu}\leq 3$ and sort the components in ascending order for visualization purposes. We then set $A^{\mu}=\psi^{-1}(1/b^{\mu})$ for each $\mu$, effectively assuming the disorder-free case with all regions in routing mode (Eq. 24). We then use the double-description algorithm to identify all verities. We summarize the results of this analysis by plotting the number of fracture variables $f$ against the log-number of vertices for many realizations of the linear program with $M\in\{5,6,7\}$ (Fig. 6(b), left). The typical number of vertices grows exponentially in $M$. $f$ is negatively correlated with the number of vertices and is at most $M-1$. With the exception of $f=3$, we find that, for a given choice of $\bm{b}$, all fracture variables correspond to the incoming and outgoing currents for a single region, specifically, the one with largest $b^{\mu}$ (Fig. 6(b), center). For $f=3$, there is an additional configuration where the variables corresponding to all three currents between the three regions with the largest values of $b^{\mu}$ are fracture variables. We visualized all $n=21$ vertices for example realizations of the linear program with $M=7$ and $f\in\{3,4,5\}$ (Fig. 6(b), right). Choices of $\bm{b}$ that lead to larger numbers of fracture variables tend to have nonuniform components (Fig. 6(c)).
We confirmed that the topology predicted by fracture variables matches that of the current-space manifold. For various samples of $\bm{b}$, we evolved the disorder-free DMFT equations with a large number of initial conditions until convergence to fixed points. We then applied t-SNE nonlinear dimensionality reduction to the collection of fixed points [53]. As expected, the number of distinct clusters was $2^{f}$ in all cases, as visualized in Fig. 6(d). Each cluster has some spread corresponding to the continuous dimensions of variation on the manifold.
@@
The dimension and topology of the current-space manifold are determined by the number of regions in routing mode and the values of $A^{\mu}$, respectively. Modifications to these quantities can be achieved by adjusting the parameters $a^{\mu}$, $b^{\mu}$, and $g^{\mu}$. Therefore, dynamically altering the strengths of interactions (for instance, through neuromodulation) offers a method to maintain an attractor manifold with a variable dimension and number of connected components, without the need to construct a completely new network architecture. This adaptability could be advantageous for responding to various cognitive tasks or environmental conditions, where the computational demands on the attractor system may change rapidly and significantly.
VI Asymmetric effective interactions
So far, we have restricted our analysis to effective interactions with a symmetry constraint. We now lift this constraint, allowing the effective-interaction tensor to have arbitrary elements. In this case, the current dynamics need not be dominated by fixed points. To analyze the dependence of the current dynamics on the geometric structure of the connectivity, we turn to the spectrum of $\hat{T}^{\mu\nu,\rho\sigma}$. When this matrix has a real leading eigenvalue (where “leading” refers to having greatest real part), the current dynamics tend to converge to fixed points. When there is instead a complex-conjugate pair of leading eigenvalues, particularly with large imaginary part, the currents typically show limit cycles. We have not found chaotic attractors in the currents.
We characterize the relationships between the current dynamics, within-region chaotic fluctuations, and the leading eigenvalue of $\hat{T}^{\mu\nu,\rho\sigma}$ through the following analysis. We consider networks of $R=2$ regions with disorder levels $g^{1}=g^{2}=1.5$. For each of many complex numbers $\lambda$ in a fixed region of the upper half-plane, we generate 50 random effective-interaction tensors $T^{\mu\nu\rho}$ whose associated matrix $\hat{T}^{\mu\nu,\rho\sigma}$ has leading eigenvalue $\lambda$ (up to the discretization of the $\lambda$ grid). We allow for nonzero within-region rank-one terms, corresponding to $T^{\mu\mu\mu}\neq 0$, but this is not essential for the results of this analysis. For each tensor, we numerically solve the nonstationary DMFT equations on a two-dimensional grid to obtain both the two-point functions $\Delta^{\mu}(t,t^{\prime})$ and currents $S^{\mu\nu}(t)$. Rather than $\Delta^{\mu}(t,t^{\prime})$ itself, we focus on the normalized quantity
$\hat{\Delta}^{\mu}(\tau)=\lim_{t\rightarrow\infty}\frac{\Delta^{\mu}(t,t+\tau)}{\sqrt{\Delta^{\mu}(t)\Delta^{\mu}(t+\tau)}}.$ (36)
To obtain the limit $t\rightarrow\infty$ in this expression, it is sufficient to evolve the mean-field dynamics past an initial transient. Chaotic fluctuations in region $\mu$ are indicated by $\hat{\Delta}^{\mu}(\tau)$ not returning to its equal-time value, ${\hat{\Delta}^{\mu}(0)=1}$, with increasing $\tau$. When $\hat{\Delta}^{\mu}(\tau)$ decays to a nonzero value, region $\mu$ displays chaotic fluctuations with underlying structure due to currents. This structure can also be seen in the currents themselves. Conversely, $\hat{\Delta}^{\mu}(\tau)$ decaying to zero indicates that there are only chaotic fluctuations in region $\mu$.
Figure 7: Portrait illustrating the dependence on the leading eigenvalue of $\hat{T}^{\mu\nu,\rho\sigma}$ of the qualitatively distinct dynamic behaviors realized in multiregion networks with asymmetric effective interactions among regions. Here, $R=2$. (a) Portrait of the behavior as a function of the leading eigenvalue. (b) Entropy of the five-way categorical distribution over dynamic behaviors at each point $\lambda$ in (a). (c) Incoming and outgoing currents (top) and two-point functions (bottom) for both regions displaying the dynamic behaviors indicated in (a).
We summarize the results of this analysis in a two-dimensional portrait, Fig. 7(a), showing the most common dynamic behavior across samples of $T^{\mu\nu\rho}$ for each $\lambda$. As the real part of $\lambda$ increases while the imaginary part remains small, the system exhibits a progression of most-common states from chaos with no currents (Fig. 7(c), “chaos”); to a state where the currents settle to a fixed point amid chaotic fluctuations of the units (Fig. 7(c), “FP + chaos”); and, for sufficiently large real part, to a state where a fixed point in the currents suppresses chaos (Fig. 7(c), “FP”). Strikingly, a parallel series of transitions occurs when the imaginary part is larger. In this case, the system transitions from chaos with no currents, as before, to a state where the currents show a limit cycle amid chaotic fluctuations of the units (Fig. 7(c), “LC + chaos”); and, finally, to a state where a limit cycle in the currents suppresses chaos (Fig. 7(c), “LC”). The case where a limit cycle in the currents coexists with chaotic fluctuations is particularly interesting, as it shows that reliable, time-dependent routing through the system can be hidden beneath noisy-looking high-dimensional fluctuations.
How useful is the leading eigenvalue of $\hat{T}^{\mu\nu,\rho\sigma}$ as a predictor of the dynamics? We quantified this by computing the entropy of the empirical distribution over the five possible dynamic states at each $\lambda$, visualized as a heatmap in Fig. 7(b). For large imaginary part of $\lambda$, as the real part is increased from zero, there is a reliable transition from chaos to a limit cycle in the currents coexisting with chaotic fluctuations of the units, with the critical real value near $\text{Re}\lambda=1.5$, the disorder variance parameter. In regions of $\lambda$ space where pure fixed points or pure limit cycles are most common, the behavior is more variable, particularly in parts of this space where different most-common states closely intermingle.
In Fig. 7(a), a dashed circle indicates the support of the bulk of the spectrum of $J^{\mu\nu}_{ij}$. For nontrivial current dynamics to occur, the leading eigenvalue must lie outside this circle; otherwise, these activity modes are swallowed by the bulk. Thus, as the disorder variance parameter increases and the circle expands, a larger-magnitude leading eigenvalue becomes necessary to obtain nontrivial current dynamics. This implies that chaotic fluctuations within regions hinder structured cross-region communication, in analogy with the non-routing mode discussed in Sec. V.3.
Figure 8: Shaping the multiregion dynamics by modulating disorder and thus signal routing. Examples 1 and 2 show two cases where introducing disorder in region 1 in a network of $R=3$ regions switches the current dynamics from fixed points to limit cycles. (a.i and a.ii) Spectra of $\hat{T}^{\mu\nu,\rho\sigma}$, which have real leading eigenvalues (top), and of the same tensor but with region 1 silenced by setting $T^{\mu 1\rho}=0$ (bottom). Silencing region 1 generates a complex-outlier pair in both cases. (b.i and b.ii) Normalized two-point functions $\hat{\Delta}^{\mu}(\tau)$ from all three regions for increasing levels of disorder in region 1, parameterized by $g^{1}$. (c.i and c.ii) Time-dependent gains $\psi^{\mu}(t)$. (d.i and d.ii) Spectra of $\psi^{\nu}(t)\hat{T}^{\mu\nu,\rho\sigma}$, with different spectra superposed across $t$.
Recall that, in the case of fixed points, switching regions between routing and non-routing modes can tune the geometry and topology of the current-space attractor manifold. We now explore the analogy of this effect in the case of asymmetric effective interactions, where dynamic attractors are possible. We study two examples of networks of $R=3$ regions where increasing the disorder variance parameter $g^{1}$ in region 1, with no disorder in the other regions, induces transitions between fixed-point and limit-cycle attractors in the currents (Fig. 8). To achieve this, in both examples, we construct the effective-interaction tensor such that $\hat{T}^{\mu\nu,\rho\sigma}$ has a real leading eigenvalue, indicative of fixed points, but when entries in $T^{\mu\nu\rho}$ corresponding to region 1 are set to zero, the leading eigenvalues become a complex-conjugate pair, as shown in Figs. 8(a.i) and (a.ii).
We increase $g^{1}$ from zero while keeping $g^{2}=g^{3}=0$ fixed and numerically solve the nonstationary DMFT equations for each value of $g^{1}$. In the first example, for small $g^{1}$, the network sits at a fixed point (Fig. 8(b.i), ${g^{1}=0.3}$). $g^{1}$ surpassing a transition value less than unity gives way to a limit-cycle attractor in the currents (Fig. 8(b.i), ${g^{1}=0.6}$). This limit cycle initially dominates chaotic fluctuations as $g^{1}$ crosses unity (Fig. 8(b.i), ${g^{1}=1.5}$). Further increasing increasing $g^{1}$ induces chaotic fluctuations of the units of region 1, and the low-dimensional communication structure prevents propagation of these high-dimensional fluctuations to other areas (Fig. 8(b.i), ${g^{1}=2.7}$). The second example initially exhibits a fixed-point attractor in the currents (Fig. 8(b.ii), ${g^{1}=0.6}$), even for ${g^{1}>1}$ due to suppression of chaos in region 1 (Fig. 8(b.ii), ${g^{1}=1.5}$). Further increasing $g^{1}$ induces chaotic fluctuations in the units of region 1 alongside a fixed point in the currents (Fig. 8(b.ii), ${g^{1}=2.1}$). For sufficiently large $g^{1}$, the system transitions to a limit cycle in the currents coexisting with chaotic fluctuations localized to the units of region 1 (Fig. 8(b.ii), ${g^{1}=2.7}$). Unlike the first example, no value of $g^{1}$ supports a pure limit cycle in the currents without chaotic fluctuations of the units.
What is the nature of the routing mechanism underlying these disorder-induced phase transitions? This mechanism does not solely involve decreasing the gain $\psi^{1}(t)$ in region 1 to remove it from the network dynamics (despite that our method of identifying appropriate effective-interaction tensors involved excluding region 1). Instead, as $g^{1}$ increases, the gains of all regions stay of order unity (Figs. 8(c.i) and (c.ii)). Increasing disorder can even increase the gain in region 1 (Fig. 8(c.i), ${g=0.3}$ vs. ${g=0.6}$). This indicates that the transitions from fixed-point to limit-cycle attractors in the currents are driven by shaping the alignment of neural activity with particular subspaces rather than traditional gain modulation. Further insight into the nature of time-dependent signal routing through the system can be obtained by studying the spectrum of $\psi^{\nu}(t)\hat{T}^{\mu\nu,\rho\sigma}$ across $t$, which controls the instantaneous linear dynamics of the currents (Figs. 8(d.i) and (d.ii)). The leading eigenvalues hover around unity as the limit cycle unfolds, indicating that the dynamics of the currents are regulated through the sequential activation of subspaces and gain adjustments.
When we allow for fully unconstrained effective interactions, even for the modest values of $R$ considered so far, the dynamics become rich and highly dependent on the specific form of the effective-interaction tensor. This raises the question of whether we can glean general insights when the number of regions is large, as is the case in real neural circuits. To investigate this, we examine a model of $R=5$ regions and asymmetric effective interactions. We randomly sample an effective-interaction tensor that yields complex-conjugate leading eigenvalues, along with several other real and complex unstable modes (Fig. 9(a)). Without disorder, this effective-interaction tensor produces an intricate limit cycle in the currents (Fig. 9(b), $g=0$). Increasing the disorder variance parameter $g$ uniformly across regions reduces the complexity of the limit cycle due to chaotic fluctuations disrupting communication between regions (Fig. 9(b), $g=1$ and $g=2$). For sufficiently large $g$, the currents vanish as disordered connectivity within regions overtakes structured communication (Fig. 9(b), $g=3$). The gradual transition from a complex to a simple limit cycle, and eventually to its absence with increasing disorder, can be understood in terms of the growing radius of the spectral bulk, which swallows more and more outlier modes linked to the dynamics of the currents (Fig. 9(a), circles). Inspection of the gains $\psi^{\mu}(t)$ reveals that, rather than vanishing, gains remain of order unity (Fig. 9(c)). However, gains exhibit less complexity across time for larger $g$. Moreover, the leading eigenvalues of the spectrum of $\psi^{\nu}(t)\hat{T}^{\mu\nu,\rho\sigma}$ hover around unity, with a diminishing number of modes crossing the stability line as $g$ increases (Fig. 9(d)).
@@ VI Asymmetric effective interactions
VII Discussion
In the cases of both fixed-point and dynamic attractors, we have shown that adjusting the effective interactions or levels of disorder can shift how signals are routed through the network by activating different subspaces. This shifting does not simply silence regions, i.e., it does not reduce the gain of a region to zero. Instead, it alters which subspaces are active, causing phase transitions where currents become zero or nonzero while the gains remain of order unity. We note that to achieve phase transitions to exactly zero currents, reciprocal effective interactions among regions are required. For instance, in a feedforward chain $1\rightarrow 2\rightarrow 3$, where region 1 sends signals to region 3 via region 2, introducing any level of disorder or low-rank connectivity in region 2 does not entirely stop communication, but does attenuate it through reducing the gain in region 2, due to the feedforward communication structure.
Neural projections between brain regions are characterized by diverse biological constraints, such as excitatory-only connections, sparsity, and layer-specificity. In our approach, we assume these projections have a low-rank structure. This assumption is inspired by experimental findings of communication subspaces and abstracts away these specific biological details to facilitate general insights.
In particular, we focused on rank-one communication subspaces with jointly Gaussian loadings. This connectivity provides a starting point for studying more complicated forms of communication between areas. For example, we can extend our rank-one connectivity model to rank-$K$ subspaces, facilitating richer, higher-dimensional communication. Crucially, maintaining the ranks of these subspaces as intensive prevents high-dimensional chaotic fluctuations from propagating between regions, preserving the modularity of the disorder-based gating mechanism. While increasing the rank increases the number dynamic variables in the mean-field picture (namely, by a factor of $K$), the Gaussian distribution over loadings restricts the complexity of their effective interactions in the DMFT. An alternative is to use a mixture-of-Gaussians distribution with $C$ components, allowing for more complex interactions [54, 55]. Together, these extensions expand the effective-interaction tensor by three indices, detailed in a tensor diagram in Fig. 10.
An additional extension to consider is the inclusion of disorder in the cross-region couplings. The connectivity in this scenario is represented as
$J^{\mu\nu}_{ij}=\chi^{\mu\nu}_{ij}+m^{\mu\nu}_{i}n^{\mu\nu}_{j},$ (37)
where $\left\langle(\chi_{ij}^{\mu\nu})^{2}\right\rangle_{\bm{J}}=G^{\mu\nu}/N$. This combines the model of Aljadeff et al. [34] with our communication-subspace model. In this modified system, the current dynamics are unchanged, but two-point function dynamics, Eq. 13, and thus the time-dependent gains, are updated to
$(1+\partial_{t})(1+\partial_{t^{\prime}})\Delta^{\mu}(t,t^{\prime})=\sum_{\nu}G^{\mu\nu}C^{\mu}(t,t^{\prime})\\ +\sum_{\nu,\rho}U^{\mu\nu\rho}H^{\mu\nu}(t)H^{\mu\rho}(t^{\prime}).$ (38)
The primary new effect is the propagation of chaotic fluctuations between regions due to the high-dimensional cross-region connectivity, captured by the coupling of $\Delta^{\mu}(t,t^{\prime})$ to other $C^{\nu}(t,t^{\prime})$ for $\nu\neq\mu$ in Eq. 38. Thus, the modularity of the disorder-based gating mechanism may not be preserved. Nevertheless, this propagation of chaotic fluctuations could shape the current dynamics in interesting ways.
We now situate our model within a line of existing research. Setting aside disorder, our model involves a blockwise low-rank coupling matrix, an embodiment of a broader idea where neurons are assigned group identities and coupling statistics are based on these identities. Another embodiment of this idea is the low-rank mixture-of-Gaussians model proposed by [54, 55], where the coupling matrix is a sum of rank-one outer products with mixture-of-Gaussians loadings. In this framework, each neuronal group corresponds to a Gaussian mixture component. Our multiregion network model, of $R$ regions, is a special case of a low-rank mixture-of-Gaussians network with rank $R^{2}$ and $R$ mixture components (Appendix A). Furthermore, as described above, multiregion networks can themselves be generalized to have rank-$K$ communication subspaces and $C$ Gaussian-mixture components. This extension can be captured by a low-rank mixture-of-Gaussians construction with $KR^{2}$ rank-one terms and $CR$ mixture components. Compared to generic networks in the low-rank mixture-of-Gaussians class, multiregion networks possess a far greater degree of structure. An interesting question is whether the inductive bias corresponding to multiregion networks is advantageous in constructing models within this class.

### v3

@@ A.6 Concrete example of the linear program construction
A.7 Characterization of the exceptional fixed point
@@
A.8 Dynamical mean-field theory (DMFT) equations for symmetric effective interactions with disorder
Assuming stationarity as described in the main text, the DMFT equations become
$\psi_{0}^{\nu}$ $=\psi(\Delta^{\nu}(0)),$ (33b)
$\frac{d^{2}\Delta^{\mu}(\tau)}{d\tau^{2}}$ $=-\partial_{\Delta^{\mu}}V^{\mu}(\Delta^{\mu};A^{\mu}),$ (33c)
$V^{\mu}(\Delta^{\mu};A^{\mu})$ $=-\frac{(\Delta^{\mu})^{2}}{2}+(g^{\mu})^{2}\Phi(\Delta^{\mu},\Delta^{\mu}(0))+A^{\mu}\Delta^{\mu},$ (33d)
$A^{\mu}$ $=\sum_{\nu}(S_{0}^{\mu\nu})^{2},$ (33e)
where we replaced ${(1+d/dt)(1+d/dt^{\prime})\rightarrow 1-d^{2}/d\tau^{2}}$ in main text Eq. 9, then integrated its rhs with respect to $\Delta^{\mu}(\tau)$. To obtain $\Phi(\Delta_{12},\Delta_{11})$, we first define $C(\Delta_{12},\Delta_{11})$ by setting $\Delta_{11}=\Delta_{12}$ in main text Eq. 10 (see also SI Eq. 23b), then integrate with respect to $\Delta_{12}$.
These equations generalize main text Eq. 15 to the case where the squared $L^{2}$-norms of the rows of the current matrix differ from the equal-time two-point functions, ${A^{\mu}\neq\Delta^{\mu}(0)}$, due to chaotic fluctuations contributing to the variance of activity in addition to the currents. As in Sompolinsky et al. [29], $\Delta^{\mu}(\tau)$ acts like a Newtonian particle in a Mexican-hat potential, $V^{\mu}(\Delta^{\mu};A^{\mu})$. The values of $\psi^{\mu}_{0}$ and thus $\Delta^{\mu}(0)$ are determined by $a^{\mu}$ and $b^{\mu}$ as in the disorder-free case, but $A^{\mu}$ is as yet undetermined. We exchange the dependence of the potential on $A^{\mu}$ for a dependence on the large-$\tau$ value of the two-point function, ${\Delta^{\mu}(\infty)=\lim_{\tau\rightarrow\infty}\Delta^{\mu}(\tau)}$, which satisfies $\left.(dV^{\mu}/d\Delta^{\mu})\right|_{\Delta^{\mu}(\infty)}=0$ because the Newtonian particle must come to rest at the top of a hill to obtain a valid decaying two-point function. This condition can be expressed as
We use this to express the potential as $V^{\mu}(\Delta^{\mu};\Delta^{\mu}(\infty))$, eliminating the dependence on $A^{\mu}$. Finally, $\Delta^{\mu}(\infty)$ is determined by energy conservation, $V^{\mu}(\Delta^{\mu}(0);\Delta^{\mu}(\infty))=V^{\mu}(\Delta^{\mu}(\infty);\Delta^{\mu}(\infty))$. $A^{\mu}$ can then be found using SI Eq. 34, shown in main text Fig. 5(a), and the full form of $\Delta^{\mu}(\tau)$ is given by integrating the Newtonian dynamics, shown in main text Fig. 5(b). A similar analysis was done by Mastrogiuseppe and Ostojic [26].
A.9 Further analysis of asymmetric effective interactions
@@
When we allow for fully unconstrained effective interactions, even for the modest values of $R$ considered so far, the dynamics become rich and highly dependent on the specific form of the effective-interaction tensor. This raises the question of whether we can glean general insights when the number of regions is large, as is the case in real neural circuits. To investigate this, we examine a model of $R=5$ regions and asymmetric effective interactions. We randomly sample an effective-interaction tensor that yields complex-conjugate leading eigenvalues, along with several other real and complex unstable modes (Fig. 9a). Without disorder, this effective-interaction tensor produces an intricate limit cycle in the currents (Fig. 9b, $g=0$). Increasing the disorder variance parameter $g$ uniformly across regions reduces the complexity of the limit cycle due to high-dimensional fluctuations disrupting communication between regions (Fig. 9b, $g=1$ and $g=2$). For sufficiently large $g$, the currents vanish as disordered connectivity within regions overtakes structured communication (Fig. 9b, $g=3$). The gradual transition from a complex to a simple limit cycle, and eventually to its absence with increasing disorder, can be understood in terms of the growing radius of the spectral bulk, which swallows more and more outlier modes linked to the dynamics of the currents (Fig. 9a, circles). Inspection of the gains $\psi^{\mu}(t)$ reveals that, rather than vanishing, gains remain of order unity (Fig. 9c). However, gains exhibit less complexity across time for larger $g$. Moreover, the leading eigenvalues of the spectrum of $\psi^{\nu}(t)\hat{T}^{\mu\nu,\rho\sigma}$ hover around unity, with a diminishing number of modes crossing the stability line as $g$ increases (Fig. 9d).
A.10 DMFT with inputs
We extend our analysis to include scalar inputs $I^{\mu}(t)$ for each region $\mu$. These inputs are supplied along a vector $\bm{v}^{\mu}$, whose components are jointly Gaussian with all other connectivity vectors in the system. The high-dimensional dynamics now become:
$\frac{dx^{\mu}_{i}(t)}{dt}=-x^{\mu}_{i}(t)+\sum_{\nu=1}^{R}\sum_{j=1}^{N}J^{\mu\nu}_{ij}\phi^{\nu}_{j}(t)+v^{\mu}_{i}I^{\mu}(t)$ (35)
Applying the DMFT analysis to this extended system yields the following expanded dynamics for the currents and two-point function:
$\frac{dS^{\mu\nu}(t)}{dt}$ $=-S^{\mu\nu}(t)+\psi^{\nu}\left(t\right)\sum_{\rho=1}^{R}T^{\mu\nu\rho}S^{\nu\rho}(t)+V^{\mu\nu}I^{\nu}(t),$ (36)
$\left(1+\frac{d}{dt}\right)\left(1+\frac{d}{dt^{\prime}}\right)\Delta^{\mu}(t,t^{\prime})$ $=(g^{\mu})^{2}C^{\mu}(t,t^{\prime})+\sum_{\nu,\rho=1}^{R}U^{\mu\nu\rho}H^{\mu\nu}(t)H^{\mu\rho}(t^{\prime})$
$+w^{\mu}I^{\mu}(t)I^{\mu}(t^{\prime})+\sum_{\nu=1}^{R}W^{\mu\nu}\left[H^{\mu\nu}(t)I^{\mu}(t^{\prime})+H^{\mu\nu}(t^{\prime})I^{\mu}(t)\right].$ (37)
Here, we introduce new parameters $V^{\mu}$, $w^{\mu}$, and $W^{\mu\nu}$, defined as averages:
$V^{\mu\nu}$ $=\left\langle n^{\mu\nu}_{i}v^{\nu}_{i}\right\rangle_{\bm{J}},$ (38)
$w^{\mu}$ $=\left\langle\left(v^{\mu}_{i}\right)^{2}\right\rangle_{\bm{J}},$ (39)
$W^{\mu\nu}$ $=\left\langle m^{\mu\nu}_{i}v^{\mu}_{i}\right\rangle_{\bm{J}}.$ (40)
To demonstrate input-driven switching, we consider a simple scenario with $R=5$ regions, using the same parameterization of connectivity as in the main text Fig. 3 (middle). We activate an input in region 1, while all other inputs remain zero. The parameters are set as follows:
$V^{\mu\nu}=0.1\times\begin{pmatrix}0&0&0&0&0\\ 1&0&0&0&0\\ 1&0&0&0&0\\ 1&0&0&0&0\\ 1&0&0&0&0\end{pmatrix},\quad w^{\mu}=1,\quad W^{\mu\nu}=0.$ (41)
Fig. 10 illustrates the effect of varying the input strength in region 1 from $I^{1}=0$ (left) to $I^{1}=2$ (right). We observe a transition in which region 1 switches from exciting itself without routing to not exciting itself and routing. We note that there is an intermediate region, appearing immediately when $I^{1}$ becomes nonzero, where the input and output currents from region 1 become nonzero, but the self-excitation of region 1 persists.
Figure 10: Input-driven switching in a 5-region network. Left: No input ($I^{1}=0$). Right: Strong input ($I^{1}=2$). The transition shows region 1 changing from self-excitation without routing to routing without self-excitation.
This example demonstrates how external inputs can modulate the routing behavior in multiregion networks, providing a mechanism for flexible, context-dependent information processing.
A.11 Extension of multiregion networks to higher-rank communication subspaces with mixture-of-Gaussians loadings.
See Fig. 11.